OpenStudy (anonymous):

Haziq drops a ball from a height of H cm above the floor.After the 1st bounce,the ball reaches a height of h_1 cm where H_1=2(H)/3.After the 2nd bounce,the ball reaches a height of H_2 cm where H_2=2(H_1)/3.The ball continue bouncing in this way until it stops. Given that H=600cm,find b) the total distance, in cm,travelled by the ball until it stops. @ganeshie8

ganeshie8 (ganeshie8):

H = 600cm = 6m Plug that in and can you list the first few terms(heights of bounces) ?

OpenStudy (anonymous):

\[600,400,266\frac{ 2 }{ 3 },177\frac{ 7 }{ 9 },118\frac{ 14 }{ 27 },...\]

ganeshie8 (ganeshie8):

Do not simplify, leave 600 as it is in the terms

ganeshie8 (ganeshie8):

\(H_0=600\) \(H_1=600(2/3)\) \(H_2=600(2/3)^2\) \(H_3=600(2/3)^3\) \(\vdots\)

ganeshie8 (ganeshie8):

Those are the heights at each step, yes ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

Now, look at what the ball is doing

ganeshie8 (ganeshie8):

It is falling freely from a height of 600cm, so it must travel a distance of 600cm before hitting the ground for the first time, yes ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

After that, it reaches a height of \(H_1\) and also it falls by the same distance

ganeshie8 (ganeshie8):

If you see, except for \(H_0\), all other heights are travelled "TWO" times. One while rising, other while falling.

ganeshie8 (ganeshie8):

remember the infinite geometric series formula ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

The distances travelled by the ball looks something like below : \(D_0=600\) \(D_1=\color{red}{2}*600(2/3)\) \(D_2=\color{red}{2}*600(2/3)^2\) \(D_3=\color{red}{2}*600(2/3)^3\) \(\vdots\)

ganeshie8 (ganeshie8):

Forget \(D_0\). Observe that the other terms form a geometric progression with \(r = 2/3\)

ganeshie8 (ganeshie8):

\(D_1=\color{red}{2}*600(2/3)\) \(D_2=\color{red}{2}*600(2/3)^2\) \(D_3=\color{red}{2}*600(2/3)^3\) \(\vdots\)

ganeshie8 (ganeshie8):

first term = \(\color{red}{2}*600(2/3)\) common ratio = \(2/3\) can you find the infinite sum using the geometric series formula ?

OpenStudy (anonymous):

\[S_\infty=\frac{ a }{ 1-r }\]\[=\frac{ 800 }{ 1-\frac{ 2 }{ 3 } }\]

ganeshie8 (ganeshie8):

Yes, simplify

OpenStudy (anonymous):

\[=2400\]

ganeshie8 (ganeshie8):

Looks good. Add the initial height \(H_0\) to that...

ganeshie8 (ganeshie8):